Can someone show me the steps on how to solve this type of problem.
(a-5)/(2)= (3a)/(2)+(a-30)/(6)

Answer by checkley77(12844)

(a-5)/(2)= (3a)/(2)+(a-30)/(6)
(a-5)/2-(3a)/2=(a-30)/6
(a-5-3a)/2=(a-30)/6
(-2a-5)/2=(a-30)/6 multiply by 6.
3(-2a-5)=a-30
-6a-15=a-30
-6a-a=-30+15
-7a=-15
a=-15/-7
a=15/7 ans.